# Learning to Orchestrate Agents in Natural Language with the Conductor

### v1

@@ 1 Introduction
Through unprecedented scale and engineering efforts, modern Large Language Models (LLMs) (anthropic2025claude-sonnet4; openai2025introducing-gpt5; openai2023gpt; team2023gemini) demonstrate the ability to solve formidably complex tasks, with performance even approaching that of top human experts (gemini_gold_medal). These remarkable latent capabilities are essentially the product of training scale and better utilization of the models themselves, with a history of research showing the importance of combining and developing the two for effective utilization liu2023pre_rebuttal; brown2020language_rebuttal. However, utilizing these latent capabilities to their full potential remains a challenge even for experienced users, with manually-designed agentic workflows being critical components of commercial AI products (AmazonQDeveloper2025; Cursor2025; MicrosoftCopilot2025) while effective prompting and self-refinement strategies are still a core focus of the current research (wei2022chain; madaan2023self). Furthermore, different models are finetuned to specialize in particular datasets and domains – with no single LM universally optimal across all tasks (Chang2024LLMEvalSurvey).
Based on these considerations, we introduce the RL Conductor: a new kind of reasoning model trained with reinforcement learning (RL) (guo2025deepseek; shao2024deepseekmath) to dynamically divide challenging problems, delegate targeted subtasks, and design communication topologies for a set of worker LLM agents. Our model is itself an LLM tasked to output a sequence of workflow steps, each defined by a natural language instruction focusing on some aspect of the overall task, the assigned agent receiving that instruction, and each agent’s visibility to the other agents as they perform their role. This formulation enables our Conductor to construct entirely flexible agentic workflows customized to each input problem, with common strategies such as prompt engineering, refinement, and even meta-prompt optimization, naturally emerging from end-to-end reward maximization.
By effectively leveraging the complementary skills of its powerful worker agents, a 7B Conductor attains state-of-the-art results on challenging reasoning benchmarks such as LiveCodeBench and GPQA Diamond (Figure 1). Our systematic evaluation shows that the performance of our Conductor goes considerably beyond both traditional self-reflection strategies with any of its workers, and also prior costly multi-agent collaboration baselines that use a much larger number of agent calls (wang2024mixture; yue2025masrouter). These findings hold well beyond the set of training tasks and across a wide range of math, coding, and natural science domains, demonstrating the potential of our model to supersede manual agentic scaffolds with a general and robust end-to-end approach.
@@ 2 Reinforcement Learning and Reasoning
Recent progress in scaling the performance of LLMs by increasing test-time compute has been driven by the reinforcement learning (RL) “reasoning” paradigm, which has established itself as a ubiquitous new stage of training large-scale open and closed source models (gpt_o1; guo2025deepseek; llama4; qwen3; comanici2025gemini). The high-level recipe, introduced by the DeepSeek R1 line of work (deepseek_res_sem; shao2024deepseekmath; guo2025deepseek), optimizes an LLM $\pi_{\theta}$, using a custom system prompt, by making it generate its own completions $o_{i}$ to a set of verifiable problems $D={(q_{1},s_{1}),\dots,(q_{N},s_{N})}$. This custom system prompt instructs the model to answer each question while providing a thinking trace before its solution attempt, placing each inside appropriate <think> and <solution> tags. The rewards $r_{i}$ for each output of the model are then determined by two conditions:
@@
Using these rewards, the model is trained with GRPO  (shao2024deepseekmath), a simple online RL algorithm. GRPO uses the LLM $\pi_{\theta}$ to generate a set of $G>1$ grouped completions $\{o_{1},\dots,o_{G}\}$ for each question $q\in D$. Then, for $\beta\geq 0$ and a KL-divergence penalty to the reference model $\mathbb{D}_{\mathrm{KL}}(\cdot\|\ \pi_{\text{ref}})$, the optimization objective is given by the KL-discounted policy maximization:
@@
using the grouped completions to compute a Monte-Carlo advantage function (sem_advantage_fn):
@@ 3.1 Framing agent coordination in natural language
While training end-to-end with the conductor reward is inherently compatible with any RL algorithm (rl_alg_ppo; rl_alg_rloo), in this work, we employ the GRPO formulation described in Section 2. Training the Conductor with this simple recipe, we observe the emergence of problem breakdowns and prompt-engineered subtasks that match the strengths of each worker, together with communication strategies that combine independent attempts with final debate rounds. As shown in Figure 3 and detailed in the following section, these Conductor behaviors lead to our model quickly surpassing each of its much larger workers, yielding state-of-the-art performance far beyond manually-designed multi-agent pipelines.
@@ 4.1 Training and evaluation setup
For our main experiments, we train a small Conductor with 7B parameters with the framework detailed in Section 3, starting from a Qwen2.5 checkpoint (hui2024qwen2). Our Conductor is tasked to devise agentic workflows of up to five steps using both proprietary frontier models, such as Gemini-2.5-Pro (comanici2025gemini), Claude-Sonnet-4 (anthropic2025claude-sonnet4), and GPT-5 (openai2025introducing-gpt5), together with open-source alternatives such as DeepSeek-R1-Distill-Qwen-32B (guo2025deepseek), Gemma3-27B-instruct (gemma3), and Qwen3-32B (qwen3). Our training dataset comprises 960 problems from four reasoning domains covering a range of math, coding, and general real-world reasoning questions, selected for their difficulty and diversity. In composing this dataset, we used the seminal competition MATH corpus (hendrycks2021measuring), the multitask language comprehension task MMLU (hendrycks2020measuring), the real-world reasoning task RLPR (yu2025rlpr), and the code generation benchmark LiveCodeBench V1 (jain2024livecodebench). We empirically find that, by relying on a powerful set of workers, our framework effectively sidesteps the canonical exploration problem faced by other small models trained with RL (rl_rlt), efficiently reaching convergence with AdamW (adamw) in only 200 GRPO iterations, a small batch size of 256 samples, and without any KL regularization.
Our evaluation focuses on assessing the Conductor’s capabilities to generalize across a set of challenging reasoning tasks both within and outside its training domain. For in-domain evaluation, we use all unseen test questions from MATH500, MMLU, RLPR, and LiveCodeBench V6. For out-of-domain evaluation, we include three unseen test tasks with GPQA Diamond (rein2024gpqa), the set of diamond difficulty questions on natural science taken from the Graduate-level Google-proof Q&A benchmark, BigCodeBench (zhuo2024bigcodebench), evaluating code generation and task automation, and AIME25 (maa_aime2025), the latest set of problems used for the American Invitational Mathematics Examination. These tasks cover a diverse set of competition and graduate-level models, with neither individual open nor closed-source model currently reigning supreme. To account for empirical stochasticity, we repeat each evaluation up to 16 times based on the number of questions in each task, reporting mean accuracy and standard errors. We provide all details on training and evaluation paradigms, including full hyperparameters, in Appendix A and E.
@@ 4.2 Elevating LLMs to a new frontier with the Conductor
We first compare the Conductor’s in-domain and out-of-domain performance with the previous best prior results obtained with different state-of-the-art open and closed source models. For each of our baselines, we report the “unconstrained”, where all models utilize maximum reasoning and output token budgets, highest recorded performance across our own re-evaluations, private implementation, and online leaderboards, potentially including proprietary, undisclosed prompting and sampling strategies. Nonetheless, as shown in Table 1, our Conductor obtains substantial improvements compared to the very best baseline in each considered task, attaining state-of-the-art performance records when evaluated both in-domain and out-of-domain. For example, as shown in Fig. 1, the performance of the Conductor at the time of writing is beyond any prior LLM on the livecodebench online leaderboard [^0] https://livecodebench.github.io/leaderboard.html, even surpassing the latest OpenAI’s O-series models (gpt_o1), which were not included in our worker pool due to their exceedingly high cost. Furthermore, we see the Conductor, through its powerful coordination strategies, is able to generate performance gains across AIME25 and GPQA-Diamond in the range of 3%, which is consistent with entire generational improvements on these challenging benchmarks, mirroring the performance jump [^1] https://www.kaggle.com/benchmarks/open-benchmarks/aime-2025222https://artificialanalysis.ai/evaluations/gpqa-diamond?models=o3%2Cgpt-5 from o3 to GPT-5, for example. We believe these results strongly validate the remarkable efficacy of our Conductor – establishing a new frontier in the capabilities of language models through collective intelligence and a new kind of powerful, adaptive agentic coordination.
@@ 4.3 Controlled large-scale evaluation
Expensive multi-turn agentic baselines. We also directly compare the RL Conductor with a broad range of multi-turn baselines and prior state-of-the-art adaptive routing strategies trained and evaluated with our same worker pool. First, we consider a robust and popular self-reflection agentic approach (selfrefine; multi_agent_debate) where for all questions, each agent is prompted up to five times to revise and improve its answers, while keeping all its previous attempts in context. Moreover, we consider four expensive multi-agent routing coordination strategies, including MASRouter (yue2025masrouter), Mixture-of-Agents (MoA) (wang2024mixture), RouterDC (chen2024routerdc), and Smoothie (guha2024smoothie). In all multi-agent baselines, we train and evaluate these models with the same set of 7 agents as our Conductor, as described in Section 4.1. These prior multi-agent strategies essentially train a router classifier to construct agentic workflows by simply selecting models and/or human-designed coordination topologies from a set of pre-specified options. We note that the expressivity of these prior strategies is inherently constrained when compared to our new framework, which places complete specification freedom in the Conductor by directly using natural language as its output medium.
@@ 5 Related Work
Reinforcement learning with tools. Reinforcement learning (RL) has become an increasingly popular paradigm for the elicitation of reasoning capabilities in LLMs (guo2025deepseek; lambert2024tulu; chu2025sft). Recent works aim to extend the RL paradigm to go beyond pure textual reasoning through tool use, enhancing capabilities in geometric reasoning, complex equation solving, code execution, search-augmented question answering, and precise computation (gehring2024rlef; feng2025retool; yu2024steptool; le2022coderl; nakano2021webgpt). For example, gehring2024rlef and le2022coderl incorporate execution feedback on code synthesis and unit tests into end-to-end RL, feng2025retool use dynamic interleaving of real-time code execution, while nakano2021webgpt equip the base model with a text-based web-browser. yu2024steptool consider multi-step tool usage through step-grained reward shaping. Our framework establishes a new extension to the tool-using RL paradigm, where powerful collaborative reasoning topologies emerge from RL by equipping the base model with workflow delegation through API calling.
Multi-agent coordination. With increasingly powerful individual LLMs, recent works aim to design topological and prompt-based scaffolds to coordinate groups of agents (du2023improving; wang2024mixture; dang2025multi; madaan2023self; yue2025masrouter). wang2024mixture and du2023improving propose carefully hand-designed scaffolds to orchestrate agents within and across successive rounds, guha2024smoothie and yue2025masrouter learn embedding spaces to map queries to agents and topologies, zhuge2024gptswarm treat collaboration as a learnable graph, and chen2024routerdc learn a router to direct queries to the single best-matched agent. Our framework differs from all existing approaches by learning powerful agent coordination strategies through pure end-to-end RL, allowing the Conductor complete freedom to learn any strategy expressible in natural language.
@@ 6 Discussion and Extensions
In this work, we introduce the Conductor, a new language model trained with reinforcement learning to push the boundaries of frontier LLMs through collective intelligence and automated prompt engineering. By dividing challenging problems, delegating targeted subtasks, and designing effective communication topologies, a 7B Conductor attains state-of-the-art performance across a diverse set of highly competitive benchmarks, going well beyond manually designed agentic pipelines and expensive multi-agent baselines. Furthermore, we demonstrate that the capabilities of our pretrained Conductor can be easily extended via finetuning, giving users the ability to specify customized agent sets and unlocking a new form of test-time scaling with recursive calling to the Conductor itself. By sharing our full code with this submission, we hope to facilitate future efforts in using language models themselves as intelligent meta-agents, flexibly harnessing the complementary capabilities of a broader set of models. To this end, an exciting, unexplored extension is to go beyond LLMs alone, introducing workers with expertise in other modalities (alphafold2; pi05), allowing the Conductor to use natural language as an expressive unifying interface and tackle increasingly ambitious human challenges in fields such as biology, robotics, and beyond.
@@ A.1 Training setup
Model, optimizer, and training regime. We use Qwen2.5-7B (hui2024qwen2) as our base model with a max completion length to 1024. We train for 200 iterations, sampling 4 questions per iteration and generating 64 rollouts per question with a temperature of 1.0. We use AdamW (adamw) as optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=0.2$ and a base learning rate of 0.000001 with cosine scheduling and a warmup ratio of 0.03. We disable reference model synchronization and set the reference model KL divergence penalty to 0.

### v2

@@ 1 Introduction
Through unprecedented scale and engineering effort, modern Large Language Models (LLMs) (Anthropic, 2025; OpenAI, 2025; 2023; Team et al., 2023) demonstrate the ability to solve formidably complex tasks, with performance even approaching that of top human experts (Luong and Lockhart, 2025). These remarkable latent capabilities are essentially the product of training scale and better utilization of the models themselves, with a history of research showing the importance of combining and developing the two for effective utilization (Liu et al., 2023; Brown et al., 2020). However, utilizing these latent capabilities to their full potential remains a challenge even for experienced users, with manually-designed agentic workflows being critical components of commercial AI products (AWS, 2025; Anysphere, 2025; Microsoft, 2025) while effective prompting and self-refinement strategies are still a core focus of the current research (Wei et al., 2022; Madaan et al., 2023a). Furthermore, different models are finetuned to specialize in particular datasets and domains – with no single LM universally optimal across all tasks (Chang et al., 2024).
Based on these considerations, we introduce the RL Conductor: a new kind of reasoning model trained with reinforcement learning (RL) (Guo et al., 2025; Shao et al., 2024) to dynamically divide challenging problems, delegate targeted subtasks, and design communication topologies for a set of worker LLM agents. Our model is itself an LLM tasked to output a sequence of workflow steps, each defined by a natural language instruction focusing on some aspect of the overall task, the assigned agent receiving that instruction, and each agent’s visibility to the other agents as they perform their role. This formulation enables our Conductor to construct entirely flexible agentic workflows customized to each input problem, with common strategies such as prompt engineering, refinement, and even meta-prompt optimization, naturally emerging from end-to-end reward maximization.
By effectively leveraging the complementary skills of its powerful worker agents, a 7B Conductor attains state-of-the-art results on challenging reasoning benchmarks such as LiveCodeBench and GPQA Diamond (Figure 1). Our systematic evaluation shows that the performance of our Conductor goes considerably beyond both traditional self-reflection strategies with any of its workers, and also prior costly multi-agent collaboration baselines that use a much larger number of agent calls (Wang et al., 2024; Yue et al., 2025). These findings hold well beyond the set of training tasks and across a wide range of math, coding, and natural science domains, demonstrating the potential of our model to supersede manual agentic scaffolds with a general and robust end-to-end approach.
@@ 2 Reinforcement Learning and Reasoning
Recent progress in scaling the performance of LLMs by increasing test-time compute has been driven by the reinforcement learning (RL) “reasoning” paradigm, which has established itself as a ubiquitous new stage of training large-scale open and closed source models (Jaech et al., 2024; Guo et al., 2025; Meta AI, 2025; Yang et al., 2025; Comanici et al., 2025). The high-level recipe, introduced by the DeepSeek R1 line of work (Wang et al., 2023; Shao et al., 2024; Guo et al., 2025), optimizes an LLM $\pi_{\theta}$, using a custom system prompt, by making it generate its own completions $o_{i}$ to a set of verifiable problems $D={(q_{1},s_{1}),\dots,(q_{N},s_{N})}$. This custom system prompt instructs the model to answer each question while providing a thinking trace before its solution attempt, placing each inside appropriate <think> and <solution> tags. The rewards $r_{i}$ for each output of the model are then determined by two conditions:
@@
Using these rewards, the model is trained with GRPO  (Shao et al., 2024), a simple online RL algorithm. GRPO uses the LLM $\pi_{\theta}$ to generate a set of $G>1$ grouped completions $\{o_{1},\dots,o_{G}\}$ for each question $q\in D$. Then, for $\beta\geq 0$ and a KL-divergence penalty to the reference model $\mathbb{D}_{\mathrm{KL}}(\cdot\|\ \pi_{\text{ref}})$, the optimization objective is given by the KL-discounted policy maximization:
@@
using the grouped completions to compute a Monte-Carlo advantage function (Sutton et al., 1999):
@@ 3.1 Framing agent coordination in natural language
While training end-to-end with the conductor reward is inherently compatible with any RL algorithm (Schulman et al., 2017; Ahmadian et al., 2024), in this work, we employ the GRPO formulation described in Section 2. Training the Conductor with this simple recipe, we observe the emergence of problem breakdowns and prompt-engineered subtasks that match the strengths of each worker, together with communication strategies that combine independent attempts with final debate rounds. As shown in Figure 3 and detailed in the following section, these Conductor behaviors lead to our model quickly surpassing each of its much larger workers, yielding state-of-the-art performance far beyond manually-designed multi-agent pipelines.
@@ 4.1 Training and evaluation setup
For our main experiments, we train a small Conductor with 7B parameters with the framework detailed in Section 3, starting from a Qwen2.5 checkpoint (Hui et al., 2024). Our Conductor is tasked to devise agentic workflows of up to five steps using both proprietary frontier models, such as Gemini-2.5-Pro (Comanici et al., 2025), Claude-Sonnet-4 (Anthropic, 2025), and GPT-5 (OpenAI, 2025), together with open-source alternatives such as DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025), Gemma3-27B-instruct (Team et al., 2025), and Qwen3-32B (Yang et al., 2025). Our training dataset comprises 960 problems from four reasoning domains covering a range of math, coding, and general real-world reasoning questions, selected for their difficulty and diversity. In composing this dataset, we used the seminal competition MATH corpus (Hendrycks et al., 2021), the multitask language comprehension task MMLU (Hendrycks et al., 2020), the real-world reasoning task RLPR (Yu et al., 2025), and the code generation benchmark LiveCodeBench V1 (Jain et al., 2024). We empirically find that, by relying on a powerful set of workers, our framework effectively sidesteps the canonical exploration problem faced by other small models trained with RL (Cetin et al., 2025), efficiently reaching convergence with AdamW (Loshchilov and Hutter, 2017) in only 200 GRPO iterations, a small batch size of 256 samples, and without any KL regularization.
Our evaluation focuses on assessing the Conductor’s capabilities to generalize across a set of challenging reasoning tasks both within and outside its training domain. For in-domain evaluation, we use all unseen test questions from MATH500, MMLU, RLPR, and LiveCodeBench V6. For out-of-domain evaluation, we include three unseen test tasks with GPQA Diamond (Rein et al., 2024), the set of diamond difficulty questions on natural science taken from the Graduate-level Google-proof Q&A benchmark, BigCodeBench (Zhuo et al., 2024), evaluating code generation and task automation, and AIME25 (Mathematical Association of America, 2025), the latest set of problems used for the American Invitational Mathematics Examination. These tasks cover a diverse set of competition and graduate-level models, with neither individual open nor closed-source model currently reigning supreme. To account for empirical stochasticity, we repeat each evaluation up to 16 times based on the number of questions in each task, reporting mean accuracy and standard errors. We provide all details on training and evaluation paradigms, including full hyperparameters, in Appendix A and E.
@@ 4.2 Elevating LLMs to a new frontier with the Conductor
We first compare the Conductor’s in-domain and out-of-domain performance with the previous best prior results obtained with different state-of-the-art open and closed source models. For each of our baselines, we report the “unconstrained”, where all models utilize maximum reasoning and output token budgets, highest recorded performance across our own re-evaluations, private implementation, and online leaderboards, potentially including proprietary, undisclosed prompting and sampling strategies. Nonetheless, as shown in Table 1, our Conductor obtains substantial improvements compared to the very best baseline in each considered task, attaining state-of-the-art performance records when evaluated both in-domain and out-of-domain. For example, as shown in Fig. 1, the performance of the Conductor at the time of writing is beyond any prior LLM on the livecodebench online leaderboard [^0] https://livecodebench.github.io/leaderboard.html, even surpassing the latest OpenAI’s O-series models (Jaech et al., 2024), which were not included in our worker pool due to their exceedingly high cost. Furthermore, we see the Conductor, through its powerful coordination strategies, is able to generate performance gains across AIME25 and GPQA-Diamond in the range of 3%, which is consistent with entire generational improvements on these challenging benchmarks, mirroring the performance jump [^1] https://www.kaggle.com/benchmarks/open-benchmarks/aime-2025222https://artificialanalysis.ai/evaluations/gpqa-diamond?models=o3%2Cgpt-5 from o3 to GPT-5, for example. We believe these results strongly validate the remarkable efficacy of our Conductor – establishing a new frontier in the capabilities of language models through collective intelligence and a new kind of powerful, adaptive agentic coordination.
@@ 4.3 Controlled large-scale evaluation
Expensive multi-turn agentic baselines. We also directly compare the RL Conductor with a broad range of multi-turn baselines and prior state-of-the-art adaptive routing strategies trained and evaluated with our same worker pool. First, we consider a robust and popular self-reflection agentic approach (Madaan et al., 2023b; Du et al., 2023b) where for all questions, each agent is prompted up to five times to revise and improve its answers, while keeping all its previous attempts in context. Moreover, we consider four expensive multi-agent routing coordination strategies, including MASRouter (Yue et al., 2025), Mixture-of-Agents (MoA) (Wang et al., 2024), RouterDC (Chen et al., 2024), and Smoothie (Guha et al., 2024). In all multi-agent baselines, we train and evaluate these models with the same set of 7 agents as our Conductor, as described in Section 4.1. These prior multi-agent strategies essentially train a router classifier to construct agentic workflows by simply selecting models and/or human-designed coordination topologies from a set of pre-specified options. We note that the expressivity of these prior strategies is inherently constrained when compared to our new framework, which places complete specification freedom in the Conductor by directly using natural language as its output medium.
@@ 5 Related Work
Reinforcement learning with tools. Reinforcement learning (RL) has become an increasingly popular paradigm for the elicitation of reasoning capabilities in LLMs (Guo et al., 2025; Lambert et al., 2024; Chu et al., 2025). Recent works aim to extend the RL paradigm to go beyond pure textual reasoning through tool use, enhancing capabilities in geometric reasoning, complex equation solving, code execution, search-augmented question answering, and precise computation (Gehring et al., 2024; Feng et al., 2025; Yu et al., 2024; Le et al., 2022; Nakano et al., 2021). For example, Gehring et al. (2024) and Le et al. (2022) incorporate execution feedback on code synthesis and unit tests into end-to-end RL, Feng et al. (2025) use dynamic interleaving of real-time code execution, while Nakano et al. (2021) equip the base model with a text-based web-browser. Yu et al. (2024) consider multi-step tool usage through step-grained reward shaping. Our framework establishes a new extension to the tool-using RL paradigm, where powerful collaborative reasoning topologies emerge from RL by equipping the base model with workflow delegation through API calling.
Multi-agent coordination. With increasingly powerful individual LLMs, recent works aim to design topological and prompt-based scaffolds to coordinate groups of agents (Du et al., 2023a; Wang et al., 2024; Dang et al., 2025; Madaan et al., 2023a; Yue et al., 2025). Wang et al. (2024) and Du et al. (2023a) propose carefully hand-designed scaffolds to orchestrate agents within and across successive rounds, Guha et al. (2024) and Yue et al. (2025) learn embedding spaces to map queries to agents and topologies, Zhuge et al. (2024) treat collaboration as a learnable graph, and Chen et al. (2024) learn a router to direct queries to the single best-matched agent. Our framework differs from all existing approaches by learning powerful agent coordination strategies through pure end-to-end RL, allowing the Conductor complete freedom to learn any strategy expressible in natural language.
@@ 6 Discussion and Extensions
In this work, we introduce the Conductor, a new language model trained with reinforcement learning to push the boundaries of frontier LLMs through collective intelligence and automated prompt engineering. By dividing challenging problems, delegating targeted subtasks, and designing effective communication topologies, a 7B Conductor attains state-of-the-art performance across a diverse set of highly competitive benchmarks, going well beyond manually designed agentic pipelines and expensive multi-agent baselines. Furthermore, we demonstrate that the capabilities of our pretrained Conductor can be easily extended via finetuning, giving users the ability to specify customized agent sets and unlocking a new form of test-time scaling with recursive calling to the Conductor itself. By sharing our full code with this submission, we hope to facilitate future efforts in using language models themselves as intelligent meta-agents, flexibly harnessing the complementary capabilities of a broader set of models. To this end, an exciting, unexplored extension is to go beyond LLMs alone, introducing workers with expertise in other modalities (Jumper et al., 2021; Intelligence et al., 2025), allowing the Conductor to use natural language as an expressive unifying interface and tackle increasingly ambitious human challenges in fields such as biology, robotics, and beyond.
@@ A.1 Training setup
Model, optimizer, and training regime. We use Qwen2.5-7B (Hui et al., 2024) as our base model with a max completion length to 1024. We train for 200 iterations, sampling 4 questions per iteration and generating 64 rollouts per question with a temperature of 1.0. We use AdamW (Loshchilov and Hutter, 2017) as optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=0.2$ and a base learning rate of 0.000001 with cosine scheduling and a warmup ratio of 0.03. We disable reference model synchronization and set the reference model KL divergence penalty to 0.